7 You wish to finance the purchase of a mini van costing $20,100 by taking out an installment loan. The loan requires a 20% down payment and equal monthly payments of $817.40 for 24 months. What is the amount of the finance charge on this loan?
Answer #1
Present value of annuity is = P*((1-(1+r)^-n)/r
Amount to be paid as equal monthly payments = $ 817.40/.
"r" is monthly rate of interets = ?
"n" is No of months = 24 months
Present value of annuity is = (20100*80%) = $ 16,080/.
16080=817.40*((1-(1+r)^-24)/r
((1-(1+r)^-24)/r = (16080/817.40)
((1-(1+r)^-24)/r = 19.67213115
((1-(1+r)^-24) = 19.67213115*r
Using trial and error, we can find "r" as below 1.49% Approx.
Finance charge on this loan = 17.88% Approx.